The distance between two point charges is increased by 10%. The force of interaction between them with be?A)increased by 10% B) decreased by 10% C) decreased by 17% D) decreased by 21%

Grade:12

1 Answers

Norman Manchu
6 years ago
F is inversely proportional to the square of distance between the charges so F1(r1)^2=F2(r2)^2If r is increased by 10% r2=1.10r and thus F2=F|1.21Percentage change in force is given by=((F2-F1)|F1 x100 which comes out to be 17%(decreased)